OpenStudy (anonymous):

x=640 solve for y y = 40 + (x - 600)*0.35

OpenStudy (rulnick):

y = 40 + (640-600)*0.35 = 40 + 40*0.35 = can you take it from here?

OpenStudy (anonymous):

yes thank you